# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

### v2

@@ 4.6 Scaling Properties: Role of Training and Inference Sequence Lengths
We study the effect of both training and inference sequence lengths on performance across tasks, i.e., scalability of seq-JEPA in terms of context length (Figure 7). We draw inspiration from recent findings in large-scale foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations. We observe that equivariance generally improves with longer training sequences (left panel), likely due to increased exposure to action-conditioned transitions. Classification performance on 3DIEBench and STL-10 (middle panel) similarly benefits. In contrast, on CIFAR100 and Tiny ImageNet with synthetic augmentations, longer training sequences slightly decrease classification performance. We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e. seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observations. In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased training sequence length. Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation.
@@ 5 Related Work
Non-Generative World Models and Joint-Embedding Predictive Architectures. Non-generative world models predict the effect of transformations or actions directly in latent space, avoiding reconstruction in pixel space. This includes contrastive SSL methods that model transformed views from context representations (van den Oord et al., 2019; Gupta et al., 2024), as well as approaches in model-based reinforcement learning (RL) to improve sample efficiency, generate intrinsic rewards, or capture environment transitions (Schwarzer et al., 2021; Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023; Guo et al., 2022). Joint-embedding predictive architectures (LeCun, 2022) form a structured subclass of non-generative world models. Rather than aligning or contrasting representations directly, they introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view. Examples include I-JEPA(Assran et al., 2023), which predicts masked regions from positional cues, and IWM (Garrido et al., 2024), which conditions on augmentation parameters. JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in structured predictive representation learning.
@@
Equivariant SSL. Equivariant SSL methods aim to retain transformation-specific information in the latent space, typically by augmenting invariant objectives with an additional equivariance term. Some approaches directly predict transformation parameters (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024; Dangovski et al., 2022). Methods such as EquiMod (Devillers and Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space via a predictor in addition to their invariant objective. SEN (Park et al., 2022) similarly predicts transformed representations but omits the invariance term. Xiao et al. (2021) use contrastive learning with separate projection heads for each augmentation, treating same-augmentation pairs as negatives. ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context using a memory module and employs a dual predictor for transformation prediction to avoid collapsing to invariance. Other approaches (Shakerinava et al., 2022; Gupta et al., 2023; Yerxa et al., 2024) do not require explicit transformation parameters but instead enforce equivariance by applying the same transformation to multiple view pairs and minimizing a distance-based loss. Finally, action-conditioned JEPAs like IWM incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives, though they still face the invariance-equivariance tradeoff depending on predictor capacity (Garrido et al., 2024).
@@
Positioning of Our Work. seq-JEPA belongs to the family of joint-embedding predictive architectures and is a non-generative world model. Unlike most equivariant SSL methods, seq-JEPA does not rely on an equivariance loss or transformation prediction objective, nor does it require view pairs with matched transformations. Instead, it introduces architectural inductive biases to JEPAs (Assran et al., 2023; Garrido et al., 2024) to learn separate invariant and equivariant representations, which emerge naturally from its action-conditioned sequential predictive learning. Furthermore, in contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al. (2019) using a transformer decoder projector conditioned on previous views via key-value caching, seq-JEPA operates on sequences of action-observation pairs in an online end-to-end manner by incorporating a sequence model (e.g. a transformer encoder) as a learned working memory during both training and inference. Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation (e.g., predictive learning across saccades and path integration).
@@ 6 Limitations and Future Perspectives
We have validated viability of seq-JEPA across a range of transformations in the image domain. Here, we discuss limitations and possible future directions. First, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization. Second, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the ResNet backbone is also trained end-to-end. The sequence scaling trends observed in the paper suggest that seq-JEPA can benefit from longer context windows over representations of pre-trained backbones as in (Pang et al., 2023; Lin et al., 2024). Third, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations). Designing group-agnostic or learned transformation models (Finzi et al., 2021) without access to transformation parameters or pairs of same transformation is an open challenge in equivariant SSL that future work may tackle. Finally, our preliminary ImageNet-1k transfer results (Appendix B.7) point to potential for broader generalization. Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Song et al. (2023) could support the development of lightweight, saliency-driven agents capable of learning efficiently from partial observations in embodied settings with a limited field of vision.
@@ A.1 Data preparation
STL10. In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE [Linardos et al., 2021], resize the output saliencies back to $96\times 96$, and store them alongside original images. We use normalization parameters $\mu[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively. After sampling fixations from saliencies, the patches that are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$). For IoR, we zero-out a circular area with radius of $16$ around each previous fixation.
@@ A.2 Training and evaluation details
Hardware. Each experiment was run on a single NVIDIA A100 GPU with 40GB of accelerator RAM. A single run of seq-JEPA on 3DIEBench with a training sequence length of three takes around 15 hours.
@@ B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
Table 5 reports performance of seq-JEPA and several equivariant baselines when conditioned on both rotation and color in 3DIEBench. All methods suffer a drop in classification performance and become highly sensitive to color. Similar performance degradations have been previously observed with 3DIEBench [Garrido et al., 2023, Gupta et al., 2024], though without a clear explanation. One plausible explanation aligned with Principle II in Wang et al. [2024], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance). Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
@@
B.2 Measuring invariance of aggregate representations
In order to have a more direct measure of invariance of the aggregate representation, we trained a model on 3DIEBench with $M_{val}=3$ and rotation conditioning, then attempted to predict the individual rotation parameters of each input view from $z_{\text{AGG}}$. To prevent leakage of rotation information during prediction, all action embeddings were zeroed out. The resulting mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; indicating that the aggregate representation contains little to no rotation-specific information.
B.3 Complete evaluation results for linear probing on top of aggregate representations
@@
B.4 Comparison of evaluation results on encoder representations and aggregate representations
@@
B.5 Additional UMAP Visualizations
@@
Figure 13: 2-D UMAP projections of seq-JEPA representations on STL-10 with action conditioning ($M_{tr}=M_{val}=4$). Encoder outputs are color-coded by class (top-left) and by $X$/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
@@
B.6 Details of Path Integration Experiments.
@@
B.7 Transfer learning results on ImageNet-1k
@@
We extract saliency maps for ImageNet-1k images using DeepGaze IIE, resize them to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this out-of-distribution (OOD) evaluation setting. Figure 15 shows top-1 linear probe accuracy on ImageNet-1k validation set across varying inference sequence lengths ($M_{val}$). For both patch sizes, performance improves with longer inference sequences, validating seq-JEPA’s ability to benefit from extended context even in this difficult OOD ImageNet-1k setting. These results echo findings in the main experiments for different training and inference sequence lengths and confirming the possibility of model’s scalability in terms of data, parameter count, and compute.

### v3

@@ 4.6 Scaling Properties: Role of Training and Inference Sequence Lengths
We study the effect of both training and inference sequence lengths on performance across tasks, i.e., scalability of seq-JEPA in terms of context length (Figure 7). We draw inspiration from recent findings in foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations.
We observe that equivariance generally improves with longer training sequences (left panel). A possible explanation for this observation is the presence of more transitions $(z_{i},a_{i},z_{i+1})$ in working memory, which means the predictor has access to a richer context $z_{\mathrm{agg}}$. This enables the predictor to more accurately approximate the transition $p(z_{i+1}\mid z_{\mathrm{agg,i}},a_{i})$. Because accurate prediction requires the model to preserve and utilize information about $a_{t}$, the encoder is implicitly encouraged to learn structured, more equivariant representations.
Classification performance on 3DIEBench and STL-10 (middle panel) benefits from a longer training sequence. In contrast, on CIFAR100 and Tiny ImageNet with synthetic augmentations, longer training sequences slightly decrease classification performance. We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e. seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observations. In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased training sequence length. Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation.
@@ 5 Related Work
Non-Generative World Models and Joint-Embedding Predictive Architectures. Non-generative world models predict the effect of transformations or actions directly in latent space, avoiding reconstruction in pixel space. This includes contrastive SSL methods that model transformed views from context representations (van den Oord et al., 2019; Gupta et al., 2024), as well as approaches in model-based reinforcement learning (RL) to improve sample efficiency, generate intrinsic rewards, or capture environment transitions (Schwarzer et al., 2021; Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023; Guo et al., 2022). Joint-embedding predictive architectures (LeCun, 2022) form a subclass of non-generative world models. They introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view. Examples include I-JEPA(Assran et al., 2023), which predicts masked regions from positional cues, and IWM (Garrido et al., 2024), which conditions on augmentation parameters. JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in representation learning and world modeling.
@@
Equivariant SSL. Equivariant SSL methods aim to retain transformation-specific information in the latent space, typically by augmenting invariant objectives with an additional equivariance term. Some approaches directly predict transformation parameters (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024; Dangovski et al., 2022). Methods such as EquiMod (Devillers and Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space via a predictor in addition to their invariant objective. SEN (Park et al., 2022) similarly predicts transformed representations but omits the invariance term. Xiao et al. (2021) use contrastive learning with separate projection heads for each augmentation, treating same-augmentation pairs as negatives. ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context and employs a dual predictor for transformation prediction to avoid collapsing to invariance. Other approaches (Shakerinava et al., 2022; Gupta et al., 2023; Yerxa et al., 2024) do not require explicit transformation parameters but instead enforce equivariance by applying the same transformation to multiple view pairs and minimizing a distance-based loss. Action-conditioned JEPAs incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives (Garrido et al., 2024). Chavhan et al. (2023) use an ensemble of heads trained on top of a pre-trained SSL encoder to span a diverse spectrum of transformation sensitivities across the latent spaces of each head. Downstream probes then learn to linearly combine these feature heads depending on the desired invariance-equivariance trade-off.
@@
Positioning of Our Work. seq-JEPA belongs to the family of joint-embedding predictive architectures and is a non-generative world model. Unlike most equivariant SSL methods, seq-JEPA does not rely on an equivariance loss or transformation prediction objective, nor does it require view pairs with matched transformations. Instead, it leverages action conditioning and architectural inductive biases to learn two separate invariant and equivariant representations. In contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al. (2019) using a transformer decoder projector conditioned on previous views via key-value caching, seq-JEPA operates on sequences of action-observation pairs in an online end-to-end manner by incorporating a sequence model (e.g. a transformer encoder) as a learned working memory during both training and inference. Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation.
@@ 6 Limitations and Future Perspectives
We have validated viability of seq-JEPA across a range of transformations in the image domain. Here, we discuss limitations and possible future directions. First, seq-JEPA is capable of autoregressive prediction in time, and therefore, can be leveraged for autoregressive latent planning in control tasks. Second, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization. Third, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the backbone is also trained end-to-end. The sequence scaling trends observed in the paper suggest that seq-JEPA can benefit from longer context windows over representations of pre-trained backbones as in (Pang et al., 2023; Lin et al., 2024). Fourth, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations). Designing group-agnostic or learned transformation models (Finzi et al., 2021) without access to transformation parameters or pairs of same transformation is an open challenge in equivariant SSL that future work may tackle. Finally, our preliminary ImageNet-1k transfer results (Appendix B.10) point to potential for broader generalization. Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Ego4D Song et al. (2023) could support the development of lightweight, saliency-driven agents capable of learning efficiently from partial observations in embodied settings with a limited field of vision.
@@ A.1 Data preparation
STL10. In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE [Linardos et al., 2021], resize the output saliencies back to $96\times 96$, and store them alongside original images. We use normalization parameters $\mu=[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively. After sampling fixations from saliencies, the patches that are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$). For IoR, we zero-out a circular area with radius of $16$ around each previous fixation.
@@ A.2 Training and evaluation details
Hardware. Each experiment was run on a single NVIDIA A100 GPU with 40GB of accelerator RAM.
Compute cost and FLOP analysis. We report pretraining compute time across methods (see Table 5). All experiments are run under the same configuration: a single A100 GPU, $\times128$ resolution, batch size 512, for 1000 epochs on 3DIEBench. seq-JEPA with a sequence length of one has similar runtime to other baselines (e.g., BYOL, SimCLR), while seq-JEPA with sequence length of three incurs a moderate increase in wall-clock time (15.1 GPU-hours) due to encoding additional action-view pairs.
Table 5: GPU-Hour comparison across methods
Method A100 pretraining GPU hours
SimCLR 11.4
BYOL 11.1
SIE 12.0
VICReg 10.9
EquiMod 11.8
seq-JEPA (train seq len = 1) 12.3
seq-JEPA (train seq len = 3) 15.1
seq-JEPA’s inference cost is primarily governed by two factors: sequence length and input resolution. Below, we report detailed FLOP counts (in Gigaflops) per forward pass of a single datapoint across a range of configurations (Note: these reflect inference-time cost).
Table 6: FLOPs for different seq-JEPA configurations (in gigaflops)
Resolution / Config Encoder only Post-aggregator
seq len = 1 seq len = 2 seq len = 3 seq len = 4 seq len = 8
Each view is $\times128$ 0.60G 0.63G 1.24G 1.85G 2.46G 4.9G
Views of $\times64$ vs. full image ($\times224$) 1.82G (224) / 0.15G (64) 0.18G 0.34G 0.51G 0.67G 1.32G
Views of $\times84$, vs. full image ($\times224$) 1.82G (224) / 0.30G (84) 0.33G 0.64G 0.95G 1.26G 2.51G
We observe that for single-view inference at $\times128$ resolution, the encoder requires only 0.60G FLOPs, which matches the compute cost of standard SSL baselines such as SimCLR and BYOL. As sequence length increases, the additional compute cost grows sub-linearly as additional views should be encoded and aggregated. For example, lengths 2 and 3—where invariant performance improves substantially—incur $\sim$1.2G and $\sim$1.85G FLOPs, respectively. The cost can be offset by reducing input resolution. In our saccade-based setup, we can feed low-resolution foveal glimpses (e.g., $\times64$ or $\times84$ crops sampled via saliency maps), which require only 0.15G–0.30G FLOPs per view. This enables the use of longer sequences at an overall compute cost comparable to full-resolution, single-view pipelines (e.g., compare the FLOPs of aggregating eight $\times64$ crops with encoding a single $\times224$ image in the second row of Table 6).
@@ A.3 Architectural details
VICRegTraj.
We use the same architecture as VICReg and add a trajectory loss (Equation 2 in Wang et al. [2024a]) with a coefficient $\lambda_{traj}=0.01$.
@@ B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
Table 7 reports performance of seq-JEPA and several equivariant baselines when conditioned on both rotation and color in 3DIEBench. All methods suffer a drop in classification performance and become highly sensitive to color. Similar performance degradations have been previously observed with 3DIEBench [Garrido et al., 2023, Gupta et al., 2024], though without a clear explanation. One plausible explanation aligned with Principle II in Wang et al. [2024b], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance). Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
@@
B.2 Results with a transformer projector instead of an MLP projector for baselines
By using a transformer projector for baselines that normally use an MLP projector, we tested whether transformer-based projectors alone account for any performance gains. We did not see any benefit from switching to transformer projectors in any of the baselines. Table 8 shows the transformer-projector results for the baselines on 3DIEBench (plus sign indicates a transformer projector).
Table 8: Comparison of baselines trained with transformer projectors on 3DIEBench.
[…]
SimCLR+ 77.92 0.32 0.48
BYOL+ 81.05 0.09 0.16
VICReg+ 75.10 0.15 0.28
VICRegTraj+ 77.82 0.24 0.36
SEN+ 81.74 0.31 0.44
Cond. BYOL+ 73.62 0.27 0.36
B.3 Out-of-distribution results
We created an OOD set for 3DIEBench to enable evaluation on unseen transformations. The original dataset samples rotations from $(-\pi/2,\pi/2)$. Our OOD test set instead uses the disjoint range $(-\pi,-\pi/2)\cup(\pi/2,\pi)$, ensuring no angular overlap with the training set. Rendering this dataset took around five hours on a single A6000 GPU. From Table 9, we see that all methods fail at OOD rotation decoding (with $R^{2}$s even reaching near 1.0-1.0). The sharp drop in $R^{2}$ values may be attributed to a domain shift in transformation geometry: while representations are well-aligned with in-distribution rotation trajectories (inside the range $(-\pi/2,\pi/2)$), they are not geometrically structured to extrapolate beyond this range to $(-\pi,-\pi/2)\cup(\pi/2,\pi)$. Despite the failure in OOD equivariance generalization, seq-JEPA exhibits a graceful degradation in classification accuracy compared to other baselines. We hypothesize that this robustness stems from its sequence-level aggregation mechanism: even when encoder representations become less equivariant under OOD transformations, aggregation across multiple views still filters out transformation-specific variability and recovers shared semantic content. Thus, seq-JEPA maintains object identity better under distribution shift, indicating that aggregation over input views supports invariance-demanding downstream tasks, even when equivariance is imperfect.
Table 9: Unseen Transformation Generalization (OOD Rotations)
[…]
SimCLR 63.86 (-17.27) -0.41 -0.201
BYOL 55.68 (-27.22) -0.25 -0.198
VICReg 61.28 (-19.20) -0.31 -0.206
SEN 60.03 (-23.40) -0.41 -0.201
SIE 60.19 (-17.30) -0.63 -0.200
EquiMoD 58.54 (-25.75) -0.45 -0.206
Cond. BYOL 57.91 (-24.70) -0.42 -0.202
VICRegTraj 62.94 (-18.32) -0.36 -0.214
seq-JEPA (1,1) 61.53 (-22.55) -0.69 -0.199
seq-JEPA (3,3) 65.03 (-21.11) -0.71 -0.211
B.4 Equivariance predictor evaluation metrics
Following the protocol in Garrido et al. [2023], we compute and report MRR, Hit@1, and Hit@5 metrics to evaluate predictor quality on the 3DIEBench validation set for seq-JEPA and two predictor-based baselines (Table 10). These results confirm that seq-JEPA achieves strong equivariance performance not only in terms of $R^{2}$, but also in top-rank retrieval metrics.
Table 10: Predictor evaluation metrics on 3DIEBench validation set.
Setting MRR H@1 H@5
SIE 0.319 0.215 0.404
EquiMod 0.136 0.037 0.186
seq-JEPA (1,1) 0.340 0.241 0.442
seq-JEPA (3,3) 0.388 0.273 0.468
B.5 Training with more compute
To examine convergence under a high-compute regime, we trained five models using 2000 epochs, $256\times 256$ resolution, and batch size 1024. Each of these experiments was run on 4 A100 GPUs (our main experiments were run on a single A100). The evaluation results (Table 11) confirm that seq-JEPA achieves a strong performance in the high-compute regime without suffering from a trade-off between invariance and equivariance. Importantly, our method achieves near-saturated performance already in the low-compute regime, indicating that it requires fewer steps for convergence and is less sensitive to input resolution and batch size than competing methods.
Table 11: Evaluation results under the high-compute regime.
Setting Top-1 Acc $R^{2}$ (rel rot) $R^{2}$ (abs rot) MRR H@1 H@5
SIE 82.652 0.721 0.764 0.411 0.287 0.490
SimCLR 85.961 0.473 0.609
EquiMod 86.833 0.492 0.625 0.154 0.048 0.201
seq-JEPA (1,1) 85.370 0.661 0.713 0.365 0.263 0.447
seq-JEPA (3,3) 87.581 0.736 0.781 0.419 0.282 0.483
B.6 Complete evaluation results for linear probing on top of aggregate representations
@@
B.7 Comparison of evaluation results on encoder representations and aggregate representations
@@
B.8 Additional UMAP Visualizations
@@
Figure 13: 2-D UMAP projections of seq-JEPA representations on STL-10 with action conditioning ($M_{tr}=M_{val}=4$). Encoder outputs are color-coded by class (top-left) and by XX/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
@@
B.9 Details of Path Integration Experiments.
@@
B.10 Transfer learning results on ImageNet-1k
@@
We extract saliency maps for ImageNet-1k images. Saliencies were extracted using DeepGaze IIE with the MIT1003 centerbias prior [Linardos et al., 2021]. Maps are saved at native resolution (matching the original ImageNet image dimensions) and normalized to probability distributions. For our transfer learning experiments we resize the images and saliencies to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this evaluation setting. Figure 15 shows top-1 linear probe accuracy on ImageNet-1k validation set across varying inference sequence lengths ($M_{val}$). For both patch sizes, performance improves with longer inference sequences, validating seq-JEPA’s ability to benefit from extended context even in this difficult OOD ImageNet-1k setting. These results echo findings in the main experiments for different training and inference sequence lengths and confirming the possibility of model’s scalability in terms of data, parameter count, and compute.